Every student in a school studies either mathematics or physics. Two-thirds of the students study mathematics and one-half of the students study physics. If seven students study both subjects, what is the total number of students in the school?
(A) 42 (B) 49 (C) 84 (D) 98 (E) None of these

Answer by checkley77(12844)

2x/3+x/2-7=x
(2*2x+3*x)/(2*3)-7=x
(4x+3x)/6-7=x
7x/6-7=x
7x/6-x=7
7x/6-6x/6=7
x/6=7
x=7*6
x=42 is the total number of students.
Proof:
2*42/3+42/2-7=42
84/3+21-7=42
28+21-7=42
42=42